) Solve the ODE y+4y+4y= e−2x using the method of variation of parameters x3


1
Expert's answer
2020-10-20T18:37:29-0400
"Solution"

Given that:-

"y''+4y'+4y=e^{-2x}"

In operator form

"(D''+4D+4)y=e^{-2x}"

Where "f(D)=(D^{2}+4D+4)"

"R=e^{-2x}"

The A.E is

"f(m)=0 \implies m^2+4m+4=0 \implies (m+2)^2=0\\
\implies m = -2, -2\\
\therefore y_c=C_1e^{-2x}+C_2xe^{-2x}\\
y_p=AU+BV\\
U=e^{-2x}\ ;\ U'=-2e^{-2x}\ ;\ \\
V=xe^{-2x}\ ;\ V'=-2xe^{-2x}+e^{-2x}\ \\
UV'+VU'=-2xe^{-4x}+e^{-4x}+2xe^{-4x}=e^{-4x}\\
\\
\therefore A=\intop \frac{-VR}{UV'-VU'}dx= - \intop \frac{xe^{-4x}}{e^{-4x}}dx = -\intop xdx\\
= -\intop xdx + \intop [\intop xdx]dx\\
= -\frac{x^2}{2}+\intop \frac{x^2}{2}dx = -\frac{x^2}{2} + \frac12 \intop x^2dx\\
\implies -\frac{x^2}{2} + \frac12 \cdot \frac{x^3}{3} = -\frac{x^2}{2} + \frac {x^3}{6}\\

\therefore B=\intop \frac{-UR}{UV'-VU'}dx= \intop \frac{e^{-2x} \cdot e^{-2x}}{e^{-4x}}dx= \intop dx\\
= \intop dx-\intop [\intop dx]dx\\
=x-\intop xdx = x-\frac{x^2}{2}\\
\implies y=y_c+y_{p}=C_1e^{-2x}+C_2xe^{-2x}+(x-\frac{x^2}{2})e^{-2x}+ (\frac{x^3}{6} +\frac{x^2}{2})e^{-2x}"